A steam pipe is covered with 1.50 cm thick insulating material with a thermal conductivity of 0.200 cal/cm·°C·s. How much energy is lost every second when the steam isat 190°C and the surrounding air is at 20.0°C? The pipe has a circumference of 800 cm and a length of 60.0 m. Neglect losses through the ends of the pipe.
Answer #1
thickness l = 1.5 cm
thermal conductivity K= 0.200 cal/cm·°C·s
temperature difference dt = 190°C -20.0°C = 170 oC
pipe has a circumference S= 800 cm
length of pipe L =60.0 m = 6000 cm
radius of the pipe r = S / (2*pi*L)
= 0.02122 cm
Area of cross section A = pi* r^ 2 = 1.414 * 10 ^-3 cm^ 2
energy is lost every second when the steam Q = KAdt / L
Q = 0.032 cal
answered by: rebel is heartbroken